# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v2

@@ 8.4. Strong cluster magnification and unique chains
Let $M/K$ be obtained by strong cluster magnification from $L/K$. Then we have that $M\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $M/K$ $\iff$ $L\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $L/K$.
@@
It is enough to show that $\tilde{M}^{N_{G^{\prime}}(H^{\prime})}=\tilde{L}^{N_{G}(H)}$ where $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$. After identifying the groups in Prop. 4.1.4, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$. Hence $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times R$. Now $\tilde{L}=\tilde{M}^{1\times R}$. Hence $\tilde{M}^{N_{G}(H)\times R}\subseteq\tilde{L}$ and ${\rm Gal}(\tilde{L}/\tilde{M}^{N_{G^{\prime}}(H^{\prime})})=N_{G}(H)$. Thus we are done. ∎
@@
Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$ as in Def 4.1.1. Then we have
@@
    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}$ is the unique ascending chain for $L/K$ $\iff$ $K\subset F_{1}F\subsetneq\dots\subsetneq F_{k}F$ is the unique ascending chain for $M/K$ for $F_{i}\subset L$ for all $i$.
@@
(1) Let $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$. After identifying the groups in Prop 4.1.4, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$. Hence ${H^{\prime}}^{G^{\prime}}=H^{G}\times 1$. Now $F=\tilde{M}^{G\times 1}$. Let $F_{1}=\tilde{L}^{H^{G}}$. Hence $F_{1}=\tilde{M}^{H^{G}\times R}$. Since $(G\times 1)\cap(H^{G}\times R)=(H^{G}\times 1)$. Hence $\tilde{M}^{H^{G}\times 1}=F_{1}F$.
(2) Since $\tilde{L}=\tilde{M}^{1\times R}$. Thus $\tilde{L}\cap F_{1}F=F_{1}$. Since $M=LF$. Hence by Thm 8.3.1 (2), we have

### v3

@@ 8.4. Strong cluster magnification and unique chains
Let $M/K$ be obtained by strong cluster magnification from $L/K$ with $r_{K}(L)\neq 1$. Then we have that $M\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $M/K$ $\iff$ $L\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $L/K$.
@@
It is enough to show that $\tilde{M}^{N_{G^{\prime}}(H^{\prime})}=\tilde{L}^{N_{G}(H)}$ where $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$. After identifying the groups in Prop. 4.1.5, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$. Hence $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times R$. Now $\tilde{L}=\tilde{M}^{1\times R}$. Hence $\tilde{M}^{N_{G}(H)\times R}\subseteq\tilde{L}$ and ${\rm Gal}(\tilde{L}/\tilde{M}^{N_{G^{\prime}}(H^{\prime})})=N_{G}(H)$. Thus we are done.
We could also see the last fact in the following way. By Lemma 2.2.2, $\tilde{L}\cap F=K\iff N\cap F=K\ \text{and}\ \tilde{L}\cap NF=N$. We also have $(\tilde{L})(NF)=\tilde{L}F=\tilde{M}$. Hence $Gal(NF/N)\cong Gal(F/K)$ and $Gal(\tilde{M}/N)=Gal(\tilde{L}/N)\times Gal(NF/N)=N_{G}(H)\times R$. ∎
Remark 8.4.2.
If $r_{K}(L)=1$ in above theorem. Then the unique descending chain for $M/K$ is $M\supsetneq L$.
@@
Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$ as in Def 4.1.1 with $t_{K}(L)\neq 1$. Then we have
@@
    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}$ is the unique ascending chain for $L/K$ $\iff$ $K\subsetneq F_{1}F\subsetneq\dots\subsetneq F_{k}F$ is the unique ascending chain for $M/K$ for $F_{i}\subset L$ for all $i$.
@@
(1) Let $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$. After identifying the groups in Prop. 4.1.5, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$. Hence ${H^{\prime}}^{G^{\prime}}=H^{G}\times 1$. Now $F=\tilde{M}^{G\times 1}$. Let $F_{1}=\tilde{L}^{H^{G}}$. Hence $F_{1}=\tilde{M}^{H^{G}\times R}$. Since $(G\times 1)\cap(H^{G}\times R)=(H^{G}\times 1)$, we get $\tilde{M}^{H^{G}\times 1}=F_{1}F$.
(2) Since $\tilde{L}=\tilde{M}^{1\times R}$. Thus $\tilde{L}\cap F_{1}F=F_{1}$. Since $M=LF$, by Thm. 8.3.1 (2), we have
@@
Remark 8.4.4.
If $t_{K}(L)=1$ in the above theorem. Then the unique ascending chain for $M/K$ is $K\subsetneq F$.
9. Properties of ascending index
Recall that the ascending index $t_{K}(L)$ of $L/K$ was defined to be the degree $[F:K]$ in Thm.  7.2.1.
Proposition 9.0.1.
If $M/L/K$ are extensions, then $t_{K}(L)|t_{K}(M)$.
Proof.
Let $F/K$ and $F^{\prime}/K$ be the unique intermediate extensions as in Thm. 7.2.1 for $L/K$ and $M/K$ respectively. Since $F^{\prime}/K$ is Galois with maximum possible degree contained in $M$, we conclude that $F\subset F^{\prime}$. Thus $[F:K]\ |\ [F^{\prime}:K]$. ∎
Remark 9.0.2.
The above proposition tells that the analogue of the notion of weak magnification always holds for ascending index.
From Lemma 8.1.6, we have the following base change result for ascending index.
Proposition 9.0.3.
Let $L/K$ be a finite extension and let $K^{\prime}/K$ be a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$. Then
$t_{K}(L)=t_{K^{\prime}}(LK^{\prime}).$
By proof of Thm. 8.4.3 (1) and Thm. 7.2.1 (2), we have the following analogue of the Cluster Magnification Theorem Thm. 2.2.3.
Theorem 9.0.4.
Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then $t_{K}(M)=d\ t_{K}(L)$ and $u_{K}(M)=u_{K}(L)$.
The following theorem is an analogue of Thm. 3.1.1.
Theorem 9.0.5.
Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
Proof.
Suppose $t=1$. By Lemmas 3.1.2 and 3.1.3 and Thm. 3.2.4 and Thm. 7.3.6 we have $L=L_{1}$ which satisfies $t_{K}(L)=1$.
Now suppose $t=n$. By Thm. 3.1.1 for $r=n$, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n$. For that $L/K$ we have $t_{K}(L)=n$.
Now suppose $1<t<n$. Hence $1<n/t<n$. By Thm. 3.1.1 for $r=n/t$ and Thm. 7.3.4, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n/t$ and $t_{K}(L)=t$.
Remark 9.0.6.
Since we have Thm. 9.0.4, we could have approached Thm. 9.0.5 in a similar way as Thm. 3.1.1 was approached in [5] using Thm. 2.2.3. This approach would have left us with some cases not covered as in Thm. 2 in [5].
The following theorem is an analogue of Thm. 2 in [5].
Theorem 9.0.7.
  1. (1)
    Let $K$ be a number field. For any integers $u\geq 3,t\geq 1$ there exists $L/K$ of degree $ut$ and $t_{K}(L)=t$.
  2. (2)
    Let $K=\mathbb{Q}$. For $t$ an even number, there exists $L/K$ of degree $2t$ and $t_{K}(L)=t$.
Proof.
(1) For $n=u$, we have $L_{1}/K$ with degree $u$ and satisfying $t_{K}(L_{1})=1$ as in first case of Thm. 9.0.5. By Lemma 2 in [5], there exists $F/K$ Galois of degree $t$ such that $\tilde{L_{1}}$ and $F$ are linearly disjoint over $K$. Hence by Thm. 9.0.4, $L=L_{1}F$ has degree $ut$ over $K$ and $t_{K}(L)=t$.
(2) Consider $P=\mathbb{Q}(\sqrt[4]{2})$ which has degree $4$ over $K=\mathbb{Q}$ and $t_{K}(P)=2$. For $t$ an even number, by Lemma 2 in [5], there exists $F/K$ Galois of degree $t/2$ such that $\tilde{P}$ and $F$ are linearly disjoint over $K$. Hence by Thm. 9.0.4, $L=PF$ has degree $2t$ over $K$ and $t_{K}(L)=t$.